OpenStudy (anonymous):

Find the domain of the given function. f(x) = square root of quantity x plus eight divided by quantity x plus four times quantity x minus six.

OpenStudy (anonymous):

\[\frac{ \sqrt{x+8} }{ x+4x-6 }\] simplify \[\frac{ \sqrt{x+8} }{ 5x-6 }\]

OpenStudy (anonymous):

supposed to be \[\sqrt{x+8}\] at the second function.

OpenStudy (anonymous):

no no its (x+4)(x-6) at the bottom

OpenStudy (anonymous):

so would the answer be x ≥ -8, x ≠ -4, x ≠ 6

OpenStudy (anonymous):

Domain is the range of x value where y is not undefined when plugging in the x Ways when y is undefined when square root is x<0 or less than zero when the denominator is zero.

OpenStudy (anonymous):

Yes, you are correct.

OpenStudy (anonymous):

the wording is weird on that one.

OpenStudy (anonymous):

it is haha and thanks!

OpenStudy (anonymous):

No problem.